Write down the Cartesian equation of a parabola with vertex at the origin and the

focus at the point (0, −2).


1
Expert's answer
2021-07-23T09:55:47-0400

Solution.

"(x-h)^2 = 4p(y-k)"

where "(h,k) - vertex", so in our case it is "(0,0)" and distance from vertex to focus is "p" = -2



"(x-0)^2 = 4*(-2)*(y-0)"


"x^2 = -8y"


"y =- 1\/8 * x^2"

Answer: "y = -1\/8 * x^2"